# Engineering impurity Bell states through coupling with a quantum bath

### v1

@@ I Introduction
Great advancements in the development of techniques to experimentally realise, control and measure various quantum systems made from single or few particles are currently fuelling a second quantum revolution [1, 2, 3]. The resulting quantum technologies rely crucially on the possibility to create and maintain entanglement in these systems [4], and one of the most useful states is the maximally entangled Bell state [5]. In this state two systems, A and B, are maximally correlated and each of them holds all possible information about some observable from the other. Therefore creating and controlling correlated quantum states, especially the Bell states, in experimentally accessible systems is of importance and will be beneficial for advancing fundamental science and quantum technologies.
@@
Cold atom systems have proven over the last two decades that they are very suitable systems to control and study single- and few-particle states [6, 7]. In particular, the long-standing problem of impurities coupled to an environment has been of immense interest ever since the availability of experimental systems in recent years [8, 9, 10, 11, 12, 13, 14, 15, 16]. Cold atomic settings allow one to access different parameter regimes and therefore simulate a wider range of interesting phenomena, which are often difficult to study with the same amount of control in condensed matter setups. Impurity physics can be conveniently discussed in terms of polarons, which are quasi-particles that are dressed by the excitations of the surrounding medium [17]. One of the most fascinating aspects of polarons is the induced interactions between them mediated by the quantum many-body environment in which they are immersed. For example, two uncorrelated impurities which are coupled to the same BEC can form a bound state (known as bipolaron) due to the bath-mediated attractive interactions [18, 19, 20, 21, 22, 23]. One of the theoretical frameworks that can successfully describe impurities in ultra-cold gases is the mapping of the BEC-impurity problem to the Fröhlich polaron Hamiltonian which describes electron-phonon interactions in condensed matter physics [24, 25, 26, 27]. For a more accurate description of impurity physics that takes the exact effects of all interactions into account, one has to go beyond the Fröhlich model [28]. While in weakly interacting gases the Gross-Pitaevskii theory [29, 30, 31] (and references therein) or the quantum Brownian motion framework [19, 16, 32, 33, 34] can be used to describe the ground state of the BEC-impurities system, in the strongly interacting regime where particle-particle correlations are prominent, the system has to be solved from first principles. The latter corresponds to accurately solving many-body systems with different intra- and inter-species interactions in the continuum, which is not a trivial task due to the large Hilbert space. However, in one dimension some exact numerical tools are available and recently the non-classical correlations in stationary and out-of-equilibrium situations have been explored using exact diagonalization methods [35, 36], or the multi-layer multi-configuration time-dependent Hartree method for mixtures of identical particles [37, 38, 39, 40]. Since the impurities can become entangled with each other due to bath induced interactions, we suggest that it is possible to control correlations between impurities by engineering the interaction with the quantum bath alone.
@@
In the following we propose a method to engineer strongly-correlated atomic states in ultracold, one-dimensional interacting atomic systems using the ability to control the inter- and intra-particle interactions via Feshbach resonances. More specifically, we consider two distinguishable impurities immersed into a third larger background component comprising of a well-defined number of bosons confined in a harmonic trap (see Fig. 1). We systematically investigate the ground state of such an interacting three-component mixture by employing the improved Exact Diagonalization scheme [41] to exactly solve the many-body Schrödinger equation describing our system. As a main result of the present work we demonstrate how to fully engineer non-classical correlations between two distinguishable impurities by tuning the impurity-impurity, and bath intraspecies interactions while the impurity-bath coupling is always repulsive. We find that for strong repulsive interactions between the impurities and the bath, the impurities phase separate to the trap edges, while the bath is localized in the trap centre. This allows to describe the impurities in a discrete spatial basis of the left and right halves of the trap and facilitates the measurement of spatial entanglement and correlation between the impurities [42]. Our results show that the impurities can form maximally entangled Bell states due to the competition between the different interaction terms, however, the correlations created between the impurities and the bath can impair Bell state formation. This depends on whether the impurities are bunched or anti-bunched, which we explain through the analysis of tripartite correlation functions. Finally, we show how the properties of the bath can be tuned to reduce these higher order correlations, allowing to create strongly entangled impurity states via bath mediated interactions.
@@
This manuscript is structured as follows: Section II presents the model and quantities of interest, while Section III discusses how impurity-impurity correlations can be induced by the bosonic bath. In Section IV we discuss how to significantly enhance the impurity-impurity correlations by engineering the bath and the conclusions and outlook are drawn in Section V.
@@ II Model and quantities of interest
The three-component mixture of interacting ultra-cold bosons we consider consists of two distinguishable atoms of species A and B, which are immersed in a background cloud of $N_{\rm{C}}$ atoms of species C. The atoms A and B are assumed to have the same mass $m_{\rm{A}}=m_{\rm{B}}=m$, while the C atoms have the mass $m_{\rm{C}}$. Since at ultra-cold temperatures the s-wave scattering processes is dominant, the interaction between any two particles can be described by a contact interaction potential [43], and we assume that the strengths of all intra- and inter-component scattering lengths can be independently adjusted using Feshbach [44] or induced-confinement resonances [45]. For numerical simplicity and experimental relevance, we also assume that all particles are trapped in a one-dimensional parabolic potential with frequency $\omega$, and hence, the Hamiltonian of this system can be written as ^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\sigma}+\hat{\mathcal{W}}_{\rm{C}}+\hat{\mathcal{W}}_{\sigma\delta}$ where
@@
Since we consider only one atom each of species A and B ($N_{\rm{A}}=N_{\rm{B}}=1$) our setting resembles a two-impurity problem. We therefore term the component C as a bosonic bath, within which the intra-species interactions are characterised by $g_{\rm{C}}$. The inter-species interactions are denoted by $g_{\sigma\delta}$ with $\sigma\neq\delta\in\{\rm{A},\rm{B},\rm{C}\}$ and describe the impurity-impurity interaction $g_{\rm{AB}}$ and the impurity-bath interactions $g_{\rm{AC}}$, $g_{\rm{BC}}$. In (quasi-)one-dimensional systems, the contact interaction potential can be modeled by a bare $\delta$-function [46]. For the sake of simplicity, we rescale all lengths, energies, and coupling strengths by harmonic oscillator units given by $\sqrt{\hbar/(m\omega)}$, $\hbar\omega$, and $\sqrt{\hbar^{3}\omega/m}$, respectively.
@@
Figure 4: Same as Fig. 3, but as a function of the mass ratio $m_{\rm{C}}/m$ for the impurity-impurity coupling strength $g_{\rm{AB}}=-3$ (blue line), $g_{\rm{AB}}=0$ (red line), and $g_{\rm{AB}}=3$ (dark-green line). The inset in panel (a) shows the bipartite mutual information between two impurities for $g_{\rm{AB}}=0$ while that in panel (b) depicts the von Neumann entropy $S_{\text{AB}}^{\text{vN}}$ as a function of $N_{\text{C}}$ for the mass ratio $m_{\rm{C}}/m=1$. Note that the horizontal dashed line in panel (d) shows $E_{\text{bp}}=0$ and the x-axis range of insets in panels (a) and (d) is the same as that of the corresponding panels. In all panels, the bath is non-interacting, $g_{\text{C}}=0$, and has $N_{\text{C}}=10$ bosons.
To structure our results, we first fix some of the parameters by choosing the impurity-bath coupling as $g_{\rm{AC}}=g_{\rm{BC}}=5$ and considering the masses of all the particles to be identical $m_{\rm{C}}/m=1$. This allows us to focus on the following three representative cases for the impurity-impurity interaction: case I where $g_{\rm{AB}}<0$, case II where $g_{\rm{AB}}=0$, and case III where $g_{\rm{AB}}>0$. In the two left columns of Fig. 2 we show the OBDMs of the A (the one for B is identical for symmetry reasons) and the C components. One can see that for all different impurity-impurity interactions the impurities localise towards the trap edges, while the particles of the bath are located in the center of the trap. The total system is therefore in a regime of phase-separation due to the large impurity-bath interactions and the size of the bath. While this describes the spatial separation of the impurities, an analysis of the diagonal of the TBDM, $\rho_{\rm{AB}}^{(2)}(x_{\rm{A}},x_{\rm{B}})$ (shown in the third column of Fig. 2), reveals that the joint probability of finding impurities A and B either side of C is strongly dependent on the impurity-impurity interaction. Specifically, the impurities anti-bunch for repulsive interactions (case III $g_{\rm{AB}}=20$) and bunch for attractive interactions (case I $g_{\rm{AB}}=-5$). Bunching is also apparent for case II, when the impurity-impurity interaction is absent, as a bound state is formed due to induced attractive impurity-impurity interactions mediated by the C component. This distinctive bound state is known as the bipolaron [18, 20, 21, 22, 23]. The TBDM in Fig. 2(g) therefore suggests the combination of bath mediated attractive interactions and suitably large repulsive impurity-bath interactions allows to create a superposition state of a bipolaron localized at each side of the C component, and that this state has a large degree of spatial entanglement. We will explore this in more detail in the next section.
@@ III Bath induced impurity-impurity correlations
The precise state in which the two impurities are in can then be determined by considering the reduced density matrix of the impurities $\rho_{\rm{LR}}$, which is a positive semi-definite Hermitian operator whose elements are constructed by integrating the TBDM $\rho^{(2)}_{\rm{AB}}(x_{\rm{A}},x_{\rm{B}},x^{\prime}_{\rm{A}},x^{\prime}_{\rm{B}})$ over the corresponding spatial regions, i.e. $|\rm{L}\rangle\in(-\infty,0)$ and $|\rm{R}\rangle\in(0,+\infty)$. In Figs. 2(d,h,l) we show the reduced density matrices for the three cases we consider. One can immediately note that Fig. 2(l) precisely exhibits the density matrix of the Bell state $|\Psi^{+}\rangle=\left(|\rm{LR}\rangle+|\rm{RL}\rangle\right)/\sqrt{2}$, which is a strong signature that the impurities are maximally entangled in case III. However, while in cases I and II the occupation of the diagonal elements $|\text{LL}\rangle\langle\text{LL}|$ and $|\text{RR}\rangle\langle\text{RR}|$ indicate that the impurities form tight bound states, the coherence terms $|\text{LL}\rangle\langle\text{RR}|$ and $|\text{RR}\rangle\langle\text{LL}|$ are significantly diminished. This implies that these states are not the maximally entangled $|\Phi^{+}\rangle=\left(|\rm{LL}\rangle+|\rm{RR}\rangle\right)/\sqrt{2}$ Bell states, and that some decoherence is being caused by the coupling to the C component. Therefore, correlations between the impurities and the matter-wave barrier have a considerable effect on impurity-impurity entanglement, which is substantially different from double-well potentials created by classical optical fields as used in most current ultra-cold atomic setups [42]. We do stress however, that the matter-wave barrier and the aforementioned induced interactions are essential for the creation of the bipolaron Bell state (case II), as classical fields alone are not sufficient to create this state.
@@
For decreasing impurity interaction $g_{\rm{AB}}$ there is a sudden transition in the concurrence as the impurities transition from being anti-bunched to being bunched. For mass-balanced systems (black lines in the following graphs) this transition is signalled by a narrow dip in the concurrence to zero in Fig. 3(a) and by a maximum in the von Neumann entropy $S^{\rm{vN}}_{\rm{AB}}$ in Fig. 3(b), the latter reflecting increased correlations between the impurities and the bath. Both cases I and II lie to the left of this transition and take significantly lower values of the concurrence and also increased correlations with the bath. This implies that there is a trade-off between the impurity-bath correlations and the impurity-impurity correlations, usually referred to as entanglement monogamy [51]. What is unclear is the stark difference between case I and III, with the latter possessing maximal concurrence while still having non-zero correlations with component C.
To understand this dichotomy we first note the differences in the structure of the bunched (cases I and II) and anti-bunched (case III) states. In the latter the impurities are spatially separated and therefore the probability to find both impurities and a bath particle at the same position is negligible. However, in the case of bunching such three particle coincidences are more likely, therefore leading to the presence of three-body correlations, and we conjecture that these higher order correlations are responsible for the decoherence of the bipolaron Bell-state. To quantify the three-body correlations we compute the tripartite mutual information (TMI) [52]
@@
Here $S^{\text{vN}}_{\sigma}$ and $S^{\text{vN}}_{\sigma\delta}$ are the respective single particle and two particle von Neumann entropies, while $S^{\text{vN}}_{\text{ABC}}$ is the three particle von Neumann entropy with one particle from each component A, B, and C. The tripartite mutual information is shown in Fig. 3(c), and as expected it vanishes when the impurities are anti-bunched, while it takes finite values when the impurities are bunched. This suggests that tripartite impurity bath correlations are indeed responsible for decohering the $|\Phi^{+}\rangle$ state.
@@
where $E_{\text{2}}$ is the ground-state energy of the total system with two impurities, $E_{\text{1}}$ denotes the ground-state energy of the system with one impurity, $E_{\text{0}}$ is the energy of the system without any impurities. This is shown in Fig. 3(d), where the binding energy takes negative values when the impurities are bound together, notably occurring for zero impurity-impurity interaction at $g_{\text{AB}}=0$. The transition between bunched and anti-bunched is signalled by the binding energy taking a constant value, occurring at the same value of $g_{\text{AB}}$ as the major changes in the concurrence, the von Neumann entropy and the mutual information. We note, however, that in the anti-bunched regime the binding energy does not vanish but instead takes a small but constant positive value, which is due to the finite size of the bath.
Figure 5: The concurrence (panel a), the von Neumann entropy (panel b) as a function of the intra-component interaction strength $g_{\rm{C}}$ for $g_{\text{AB}}=-3$ (blue line), $g_{\text{AB}}=0$ (red line), and $g_{\text{AB}}=3$ (dark-green line). In all panels the mass ratio is $m_{\rm{C}}/m=1$.
Figure 6: The reduced density matrix $\rho_{\text{LR}}$ for (a) $m_{\rm{C}}/m=15;g_{\text{C}}=0$, (b) $m_{\rm{C}}/m=1;g_{\text{C}}=-5$, and $m_{\rm{C}}/m=15;g_{\text{C}}=-5$. In all panels, the impurity-impurity coupling strength is $g_{\text{AB}}=-3$, and the bath has $N_{\text{C}}=10$ bosons.

### v2

@@ I Introduction
Great advancements in the development of techniques to experimentally realise, control and measure various quantum systems made from single or few particles are currently fueling a second quantum revolution [1, 2, 3]. The resulting quantum technologies rely crucially on the possibility to create and maintain entanglement in these systems [4], and one of the most useful states is the maximally entangled Bell state [5]. In this state two systems, A and B, are maximally correlated and each of them holds all possible information about some observable from the other. Therefore creating and controlling correlated quantum states, especially the Bell states, in experimentally accessible systems is of importance and will be beneficial for advancing fundamental science and quantum technologies.
@@
Cold atom systems have proven over the last two decades that they are very suitable systems to control and study single- and few-particle states [6, 7]. In particular, the long-standing problem of impurities coupled to an environment has been of immense interest ever since the availability of experimental systems in recent years [8, 9, 10, 11, 12, 13, 14, 15, 16]. Cold atomic settings allow one to access different parameter regimes and therefore simulate a wider range of interesting phenomena, which are often difficult to study with the same amount of control in condensed matter setups. Impurity physics can be conveniently discussed in terms of polarons, which are quasi-particles that are dressed by the excitations of the surrounding medium [17]. One of the most fascinating aspects of polarons is the induced interactions between them mediated by the quantum many-body environment in which they are immersed (see Ref. [18] for a recent review). For example, two uncorrelated impurities which are coupled to the same Bose-Einstein condensate (BEC) can form a bound state (known as a bipolaron) due to the bath-mediated attractive interactions [19, 20, 21, 22, 23, 24]. One of the theoretical frameworks that can successfully describe impurities in ultra-cold gases is the mapping of the BEC-impurity problem to the Fröhlich polaron Hamiltonian which describes electron-phonon interactions in condensed matter physics [25, 26, 27, 28]. For a more accurate description of impurity physics that takes the exact effects of all interactions into account, one has to go beyond the Fröhlich model [29]. While in weakly interacting gases the Gross-Pitaevskii theory (see Refs. [30, 31, 32] and references therein) or the quantum Brownian motion framework [20, 16, 33, 34, 35] can be used to describe the ground state of the BEC-impurities system, in the strongly interacting regime where particle-particle correlations are prominent, the system has to be solved from first principles. It is also important to note that two recent works have shown that the impurity-bath correlations play a crucial role and result in the suppression of impurity self-localization [36, 37], which is not captured by the mean-field approach. This demonstrates that a full quantum treatment of the BEC-impurity problem is necessary, although it is not a trivial task to accurately solve many-body systems with different intra- and inter-species interactions in the continuum due to the exponential growth of Hilbert space. However, in one dimension some exact numerical tools are available and recently the non-classical correlations in stationary and out-of-equilibrium situations have been explored using exact diagonalization methods [38, 39], or the multi-layer multi-configuration time-dependent Hartree method for mixtures of identical particles [40, 41, 42, 43]. Since the impurities can become entangled with each other due to bath induced interactions, we suggest that it is possible to control correlations between impurities by engineering the interaction with the quantum bath alone.
@@
In the following we propose a method to engineer strongly-correlated atomic states in ultracold, one-dimensional (1D) interacting atomic systems using the ability to control the inter- and intra-particle interactions via Feshbach resonances. More specifically, we consider two distinguishable impurities immersed into a third larger background component composed of a well-defined number of bosons confined in a harmonic trap (see Fig. 1). We systematically investigate the ground state of such an interacting three-component mixture by employing the improved Exact Diagonalization scheme [44] to exactly solve the many-body Schrödinger equation describing our system (see also Appendix A). As a main result of the present paper we demonstrate how to fully engineer non-classical correlations between two distinguishable impurities by tuning the impurity-impurity, and bath intraspecies interactions while the impurity-bath coupling is always repulsive. We find that for strong repulsive interactions between the impurities and the bath, the impurities phase separate to the trap edges, while the bath is localized in the trap center. This allows one to describe the impurities in a discrete spatial basis of the left and right halves of the trap and facilitates the measurement of spatial entanglement and correlation between the impurities [45]. Our results show that the impurities can form maximally entangled Bell states due to the competition between the different interaction terms, however, the correlations created between the impurities and the bath can impair Bell state formation. This depends on whether the impurities are bunched or anti-bunched, which we explain through the analysis of tripartite correlation functions. Finally, we show how the properties of the bath can be tuned to reduce these higher order correlations, allowing one to create strongly entangled impurity states via bath mediated interactions.
@@
This paper is structured as follows: Section II presents the model and quantities of interest, while Section III discusses how impurity-impurity correlations can be induced by the bosonic bath. In Section IV we discuss how to significantly enhance the impurity-impurity correlations by engineering the bath and the conclusions and outlook are drawn in Section V.
@@ II Model and quantities of interest
The three-component mixture of interacting ultra-cold bosons we consider consists of two distinguishable atoms of species A and B, which are immersed in a background cloud of $N_{\rm{C}}$ atoms of species C. The atoms A and B are assumed to have the same mass $m_{\rm{A}}=m_{\rm{B}}=m$, while the C atoms have the mass $m_{\rm{C}}$. Since at ultra-cold temperatures the s-wave scattering process is dominant, the interaction between any two particles can be described by a contact interaction potential [46], and we assume that the strengths of all intra- and inter-component scattering lengths can be independently adjusted using Feshbach [47] or induced-confinement resonances [48]. For numerical simplicity and experimental relevance, we also assume that all particles are trapped in a one-dimensional parabolic potential with frequency $\omega$, and hence, the Hamiltonian of this system can be written as ^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\sigma}+\hat{\mathcal{W}}_{\rm{C}}+\hat{\mathcal{W}}_{\sigma\delta}$ where
@@
Since we consider only one atom each of species A and B ($N_{\rm{A}}=N_{\rm{B}}=1$) our setting resembles a two-impurity problem. We therefore term the component C as a bosonic bath, within which the intra-species interactions are characterized by $g_{\rm{C}}$. The inter-species interactions are denoted by $g_{\sigma\delta}$ with $\sigma\neq\delta\in\{\rm{A},\rm{B},\rm{C}\}$ and describe the impurity-impurity interaction $g_{\rm{AB}}$ and the impurity-bath interactions $g_{\rm{AC}}$ and $g_{\rm{BC}}$. In (quasi-)one-dimensional systems, the contact interaction potential can be modeled by a bare $\delta$-function [49]. For the sake of simplicity, we rescale all lengths, energies, and coupling strengths by harmonic oscillator units given by $\sqrt{\hbar/(m\omega)}$, $\hbar\omega$, and $\sqrt{\hbar^{3}\omega/m}$, respectively.
@@
Figure 4: Same as Fig. 3, but as a function of the mass ratio $m_{\rm{C}}/m$ for the impurity-impurity coupling strength $g_{\rm{AB}}=-3$ (blue line), $g_{\rm{AB}}=0$ (red line), and $g_{\rm{AB}}=3$ (dark-green line). The insets in panel (a) and (b) show the concurrence $\mathcal{C}$ and the von Neumann entropy $S_{\text{AB}}^{\text{vN}}$ as a function of $N_{\text{C}}$ for the mass ratio $m_{\rm{C}}/m=1$, respectively. Meanwhile, the inset in panel (c) depicts the bipartite mutual information between two impurities as a function of the mass ratio for $g_{\rm{AB}}=0$. Note that the horizontal dashed line in panel (d) shows $E_{\text{bp}}=0$ and the x-axis range of insets in panels (c) and (d) is the same as that of the corresponding panels. In all panels, the bath is non-interacting, $g_{\text{C}}=0$, and has $N_{\text{C}}=10$ bosons.
To structure our results, we first fix some of the parameters by choosing the impurity-bath coupling as $g_{\rm{AC}}=g_{\rm{BC}}=5$ and considering the masses of all the particles to be identical $m_{\rm{C}}/m=1$. This allows us to focus on the following three representative cases for the impurity-impurity interaction: case I where $g_{\rm{AB}}<0$, case II where $g_{\rm{AB}}=0$, and case III where $g_{\rm{AB}}>0$. In the two left columns of Fig. 2 we show the OBDMs of the A (the one for B is identical for symmetry reasons) and the C components. One can see that for all different impurity-impurity interactions the impurities localize towards the trap edges, while the particles of the bath are located in the center of the trap. The total system is therefore in a regime of phase-separation due to the large impurity-bath interactions and the size of the bath. While this describes the spatial separation of the impurities, an analysis of the diagonal of the TBDM, $\rho_{\rm{AB}}^{(2)}(x_{\rm{A}},x_{\rm{B}})$ (shown in the third column of Fig. 2), reveals that the joint probability of finding impurities A and B on either side of C is strongly dependent on the impurity-impurity interaction. Specifically, the impurities anti-bunch for repulsive interactions (case III $g_{\rm{AB}}=20$) and bunch for attractive interactions (case I $g_{\rm{AB}}=-5$). Bunching is also apparent for case II, when the impurity-impurity interaction is absent, as a bound state is formed due to induced attractive impurity-impurity interactions mediated by the C component. This distinctive bound state is known as the bipolaron [19, 21, 22, 23, 24]. The TBDM in Fig. 2(g) therefore suggests the combination of bath mediated attractive interactions and suitably large repulsive impurity-bath interactions allows one to create a superposition state of a bipolaron localized at each side of the C component, and that this state has a large degree of spatial entanglement. We will explore this in more detail in the next section.
@@ III Bath induced impurity-impurity correlations
The precise state which the two impurities are in can then be determined by considering the reduced density matrix of the impurities $\rho_{\rm{LR}}$, which is a positive semi-definite Hermitian operator whose elements are constructed by integrating the TBDM $\rho^{(2)}_{\rm{AB}}(x_{\rm{A}},x_{\rm{B}},x^{\prime}_{\rm{A}},x^{\prime}_{\rm{B}})$ over the corresponding spatial regions, i.e. $|\rm{L}\rangle\in(-\infty,0)$ and $|\rm{R}\rangle\in(0,+\infty)$. In Figs. 2(d,h,l) we show the reduced density matrices for the three cases we consider. One can immediately note that Fig. 2(l) precisely exhibits the density matrix of the Bell state $|\Psi^{+}\rangle=\left(|\rm{LR}\rangle+|\rm{RL}\rangle\right)/\sqrt{2}$, which is a strong signature that the impurities are maximally entangled in case III. However, while in cases I and II the occupation of the diagonal elements $|\text{LL}\rangle\langle\text{LL}|$ and $|\text{RR}\rangle\langle\text{RR}|$ indicates that the impurities form tight bound states, the coherence terms $|\text{LL}\rangle\langle\text{RR}|$ and $|\text{RR}\rangle\langle\text{LL}|$ are significantly diminished. This implies that these states are not the maximally entangled $|\Phi^{+}\rangle=\left(|\rm{LL}\rangle+|\rm{RR}\rangle\right)/\sqrt{2}$ Bell states, and that some decoherence is being caused by the coupling to the C component. Therefore, correlations between the impurities and the matter-wave barrier have a considerable effect on impurity-impurity entanglement, which is substantially different from double-well potentials created by classical optical fields as used in most current ultra-cold atomic setups [45]. We do stress however, that the matter-wave barrier and the aforementioned induced interactions are essential for the creation of the bipolaron Bell state (case II), as classical fields alone are not sufficient to create this state.
@@
For decreasing impurity interaction $g_{\rm{AB}}$ there is a sudden transition in the concurrence as the impurities transition from being anti-bunched to being bunched. For mass-balanced systems (black lines in the following graphs) this transition is signaled by a narrow dip in the concurrence to zero in Fig. 3(a) and by a maximum in the von Neumann entropy $S^{\rm{vN}}_{\rm{AB}}$ in Fig. 3(b), the latter reflecting increased correlations between the impurities and the bath. Both cases I and II lie to the left of this transition and take significantly lower values of the concurrence and also increased correlations with the bath. This implies that there is a trade-off between the impurity-bath correlations and the impurity-impurity correlations, usually referred to as entanglement monogamy [54]. What is unclear is the stark difference between case I and III, with the latter possessing maximal concurrence while still having non-zero correlations with component C.
To understand this dichotomy we first note the differences in the structure of the bunched (cases I and II) and anti-bunched (case III) states. In the latter the impurities are spatially separated and therefore the probability to find both impurities and a bath particle at the same position is negligible. However, in the case of bunching such three-particle coincidences are more likely, therefore leading to the presence of three-body correlations, and we conjecture that these higher order correlations are responsible for the decoherence of the bipolaron Bell-state. To quantify the three-body correlations we compute the tripartite mutual information (TMI) [55]
@@
Here $S^{\text{vN}}_{\sigma}$ and $S^{\text{vN}}_{\sigma\delta}$ are the respective single-particle and two-particle von Neumann entropies, while $S^{\text{vN}}_{\text{ABC}}$ is the three-particle von Neumann entropy with one particle from each component A, B, and C. The tripartite mutual information is shown in Fig. 3(c), and as expected it vanishes when the impurities are anti-bunched, while it takes finite values when the impurities are bunched. This suggests that tripartite impurity-bath correlations are indeed responsible for decohering the $|\Phi^{+}\rangle$ state.
@@
where $E_{\text{2}}$ is the ground-state energy of the total system with two impurities, $E_{\text{1}}$ denotes the ground-state energy of the system with one impurity, and $E_{\text{0}}$ is the energy of the system without any impurities. This is shown in Fig. 3(d), where the binding energy takes negative values when the impurities are bound together, notably occurring for zero impurity-impurity interaction at $g_{\text{AB}}=0$. The transition between bunched and anti-bunched is signalled by the binding energy taking a constant value, occurring at the same value of $g_{\text{AB}}$ as the major changes in the concurrence, the von Neumann entropy and the mutual information. We note, however, that in the anti-bunched regime the binding energy does not vanish but instead takes a small but constant positive value, which is due to the finite size of the bath.
Figure 5: The concurrence (panel a) and the von Neumann entropy (panel b) as a function of the intra-component interaction strength $g_{\rm{C}}$ for $g_{\text{AB}}=-3$ (blue line), $g_{\text{AB}}=0$ (red line), and $g_{\text{AB}}=3$ (dark-green line). In all panels the mass ratio is $m_{\rm{C}}/m=1$.
Figure 6: The reduced density matrix $\rho_{\text{LR}}$ for (a) $m_{\rm{C}}/m=15,g_{\text{C}}=0$; (b) $m_{\rm{C}}/m=1,g_{\text{C}}=-5$; and $m_{\rm{C}}/m=15,g_{\text{C}}=-5$. In all panels, the impurity-impurity coupling strength is $g_{\text{AB}}=-3$, and the bath has $N_{\text{C}}=10$ bosons.